OpenStudy (cucacula):

Sin Theta/1+Cos Theta + 1+Cos Theta/Sin Theta = 2 Cosec Theta . prove the identity.

OpenStudy (anonymous):

Proving it will take quite a bit of work. You can verify it using other, more fundamental, identities. Looks like you want to verify that \[\large \frac{sin}{1+cos}+\frac{1+cos}{sin}=2csc\]

OpenStudy (anonymous):

Try adding the fractions on the left (remember, you'll need a common denominator), then simplify that.

OpenStudy (cucacula):

@CliffSedge can write the 1st step, i'm really don't understand ..

OpenStudy (anonymous):

I like the "bow-tie" method for adding fractions. \[\large \frac{a}{b}+\frac{c}{d}=\frac{ad+bc}{bd}\]

OpenStudy (anonymous):

a=d=sin(Θ) b=c=1+cos(Θ)

OpenStudy (cucacula):

if bd = ? @CliffSedge

OpenStudy (anonymous):

bd=sin(1+cos)

OpenStudy (anonymous):

\[\large \frac{sin}{1+cos}+\frac{1+cos}{sin}=\frac{sin^2+(1+cos)^2}{sin(1+cos)}\]

OpenStudy (anonymous):

Simplify the top and you'll see what can cancel in the bottom.

OpenStudy (cucacula):

alright i got already . thanks :) @CliffSedge im trying to do first ..